OpenStudy (anonymous):

Simplify (7t^2+9)+(6t^2+8)

OpenStudy (geneticrockhopper247):

All that you really have to do here is combine like terms: 7+6=13 (the exponent is the same, so we don't have to worry about that) and 9+8=17. So, we have 13t^2+17 as our answer. Does that make sense??